# Robustness of topological phases on aperiodic lattices

### v1

@@ 1. Introduction
We begin with some background around it. Topological insulators are materials that are insulating in the bulk but nevertheless possess metallic edges. The current flowing on the edge is usually quite robust under disorder coming from impurity of the crystal. Starting from Bellissard, van Elst and Schulz-Baldes [Bellissard-van_Elst-Schulz_Baldes:NCG_of_QHE, Bellissard:K-theory_in_solid_state_physics], the non-commutative and C*-algebraic tools have been intensively applied to the study of such materials and in particular, leading to an interpretation of the Kubo formula in the integer quantum Hall effect (IQHE) as a quantised pairing between K-theory and cyclic cohomology.
@@
These suggest to use more general point sets in $\mathbb{R}^{d}$, modelling the atomic sites of this material, as their underlying geometric spaces. The dynamics thereon are studied by more general actions other than a $\mathbb{Z}^{d}$-translation. It was explained in [Bellissard-Herrmann-Zarrouati:Hulls_aperiodic_solids] that such point sets should be Delone sets (Definition 2.14). Let $\Lambda$ be a Delone set, whose points are considered as the sites in a physical system. How should one model the observable C*-algebra $A$ from it?
The choice of an observable C*-algebra should be aligned with the following principle: it should be large enough to contain all possible Hamiltonians, and small enough to supply useful homotopy theory (K-theory). In recent years, there have been two main approaches to this modelling problem, which provide toolkits to compute numerical invariants of topological phases:
@@
  • “universal” approach describes a C*-algebra which is stable under all short-range, locally–finite-rank perturbations. This leads to a Roe C*-algebra.
The groupoid approach gives an étale groupoid $\mathcal{G}_{\Lambda}\rightrightarrows\Omega_{0}$ from a Delone set $\Lambda$, and yields a “tight-binding” groupoid C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ , cf. [Bellissard-Herrmann-Zarrouati:Hulls_aperiodic_solids, Bourne-Prodan:Noncommutative_Chern_numbers, Bourne-Mesland:Aperiodic, Mesland-Prodan:Interacting]. This is a generalisation of the well-studied periodic model: if $\Lambda=\mathbb{Z}^{d}$, then $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\simeq\mathrm{C}^{*}(\mathbb{Z}^{d})$ is the group C*-algebra of $\mathbb{Z}^{d}$. Here tight-binding means the following: the regular representation of $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ is given by a Hilbert $\mathrm{C}(\Omega_{0})$-module, or equivalently, a continuous field of Hilbert spaces over $\Omega_{0}$, where $\Omega_{0}$ is unit space of the groupoid $\mathcal{G}_{\Lambda}$. The fibres of this continuous field are canonically identified with $\ell^{2}(\omega)$, where $\omega\in\Omega_{0}$ is a Delone set, either as a translated copy of $\Lambda$, or as a weak*-limit of such sets. Thus $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ consists of families of model Hamiltonians $(H_{\omega})_{\omega\in\Omega_{0}}$, where $H_{\omega}$ acts on $\ell^{2}(\omega)$, in a covariant way with respect to the groupoid action.
The coarse-geometric approach describes a C*-algebra which is stable under all possible short-range perturbations. This gives rise to a (uniform or non-uniform) Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ from $\Lambda$, cf. [Kubota:Controlled_topological_phases, Ewert-Meyer:Coarse_geometry]. We choose the non-uniform Roe C*-algebras, whose advantages over the uniform ones were explained in [Ewert-Meyer:Coarse_geometry].
@@
Topological phases on materials are represented by a K-theory class of the C*-algebra of observables, constructed from the Hamiltonian $H$. If the system consists of conjugate-linear symmetries, then the C*-algebra is replaced by a $\mathbb{Z}/2$-graded “real” C*-algebra of the form $A_{\mathbb{C}}\otimes\mathbb{C}\ell_{p,q}$, or its corresponding real subalgebra $A_{\mathbb{R}}\otimes\mathrm{C}\ell_{p,q}$. We wish to describe the topological phases with their numerical invariants. That is, we seek for maps
(1.1) $\operatorname{KK}(\mathbb{R},A\otimes\mathrm{C}\ell_{p,q})\to\mathbb{R}$
generating a real number from a prescribed topological phase. Moreover, we wish to understand the robustness of such numbers.
@@
If we model a topological insulator on a Delone set $\Lambda$ by the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$, then its topological phase is robust in a very strong sense: such topological phases (and hence their numerical invariants as in (1.1)) are robust under any short-range, locally–finite-rank perturbation that does not close the spectral gap (and preserves the symmetry). It follows from the K-theory of (real) Roe C*-algebra (cf. (3.21)) that such a topological phase can be described completely by a $\mathbb{Z}$ or $\mathbb{Z}/2$-valued index.
@@
The groupoid model $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, on the other hand, has more involved K-theory and numerical invariants. One natural question is in which sense these topological phases are still robust. Moreover, do they lead to interesting invariants that do not occur in the coarse-geometric model? Questions related to robustness can be answered by showing that the numerical index is still continuous, even for some more general perturbations of the Hamiltonian (which does not have to belong to the observable C*-algebra), together with a “quantisation” statement that the range of the numerical index is $\mathbb{Z}$. This implies that such perturbations leave the numerical index unchanged.
@@
We describe a different approach to this robustness question by comparing the groupoid model with the coarse-geometric model on $\Lambda$ using the regular representation. The numerical invariants defined in [Bourne-Mesland:Aperiodic] factor through this representation. Therefore, since $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ is “robust”, one only needs to understand which topological phases in $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ still survive there. To this end, we introduce a handy definition of position spectral triples (Definition 2.16). Both the groupoid model and the coarse-geometric model possess such type of spectral triples. In particular, the position spectral triple $\xi^{\mathrm{Roe}}_{\omega}$ over the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ induces an isomorphism $\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))\to\mathbb{Z}$ (Theorem 3.32). Then we show that those numerical invariants of the groupoid model, given by localising the “bulk cycle” $\prescript{}{d}{\lambda}_{\Omega_{0}}$ at $\omega\in\Omega_{0}$, are pullbacks of the class of $\xi^{\mathrm{Roe}}_{\omega}$ in K-homology under the comparison *-homomorphism $\mathbb{M}_{N}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ (Theorem 4.1). Such invariants are therefore strong in the sense of [Ewert-Meyer:Coarse_geometry].
The comparison *-homomorphisms $\pi_{\omega}^{N}\colon\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ are (localised) regular representation on the tight-binding Hilbert spaces $\ell^{2}(\omega,\mathcal{K})$, where $\mathcal{K}$ is an auxilliary separable Hilbert space, considered as a generalised version of the “fundamental domain”. Such a representation, together with the position operators thereon, are zipped into the data of position spectral triples, which is compatible with the numerical invariants constructed in [Bourne-Mesland:Aperiodic, Ewert-Meyer:Coarse_geometry]. On the other hand, Kubota [Kubota:Controlled_topological_phases] has described a way to describe a comparison *-homomorphism $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\to\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ whenever $\Lambda$ is aperiodic. We explain in Example 4.4 that Kubota’s construction can be concretely realised as $\pi_{\Lambda}$; in particular, this works for any $\omega\in\Omega_{0}$ disregarding whether or not the corresponding Delone set $\omega$ is aperiodic.
@@
As another application, we show in Theorem 4.11 that certain numerical invariants on Delone sets, coming from “stacking” lower-dimensional topological phases, must be weak, i.e. unstable under perturbation. This generalises a result in [Ewert-Meyer:Coarse_geometry], which explains why certain topological phases of the periodic model ought to be weak. Ewert and Meyer compare the periodic model and the coarse-geometric model using an injective *-homomorphism $\mathrm{C}^{*}(\mathbb{Z}^{d})\hookrightarrow\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$. It vanishes on all but one $\mathbb{Z}$-component of K-theory. In particular, topological phases coming from “stacking” lower-dimensional topological phases will all vanish in $\operatorname{K}_{*}(\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))$. This allows to conclude that the induced maps $\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d}))\to\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d+1}))$ also vanishes there. We show that one can replace $\mathbb{Z}^{d+1}$ by a product Delone set of the form $\Lambda\times L$, where $\Lambda\subseteq\mathbb{R}^{m}$ and $L\subseteq\mathbb{R}^{n}$ are both Delone sets. Using the same strategy in [Ewert-Meyer:Coarse_geometry], we show that stacked topological phases factor through the Roe C*-algebra of a flasque space, which has vanishing K-theory. This allows to conclude that such topological phases and their numerical invariants have to vanish in the target Roe C*-algebra.
1.1. Future work
We remark here with some results that will appear in a forthcoming work [Li:Coarse_groupoid_robustness]. Bourne and Mesland [Bourne-Mesland:Aperiodic] have provided another way to construct numerical invariants for the groupoid model of topological phases, by composing the bulk cycle with a semi-finite spectral triple. Unlike the case studied in this present article, such numerical invariants have real numbers as their ranges, and take more than one Delone set $\omega\in\Omega$ as input. We shall show that such numerical invariants are still robust, by using a similar construction of Kubota, but with a family of coarse groupoids.
An analog of Kubota’s construction in the continuum setting is given in [Ewert-Meyer:Coarse_geometry], where the observable C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ is replaced by $\mathrm{C}^{*}(\Omega_{\Lambda})\rtimes\mathbb{R}^{d}$; and $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ is replaced by $\mathbb{R}^{d}$. Such constructions are compatible with Kubota’s construction, but one must carefully define the maps between the corresponding “tight-binding” and “continuum” versions of observable C*-algebras.
1.2. Acknowledgement
The author thanks Guo Chuan Thiang, Matthias Ludewig, Matthias Frerichs for helpful discussions and comments. In particular, the author thanks his supervisor Bram Mesland for heartfelt encouragement and a careful proofreading, helping the author to eliminate various errors in an earlier version of this work. The work is part of NWO project 613.009.142 “Noncommutative index theory of discrete dynamical systems”.
@@ 1.3. Notation and conventions
We list a few symbols and conventions that will be used in the article.
@@
Let $\Lambda$ be a discrete set. We write $\lvert x\rangle$ for the function in $\ell^{2}(\Lambda)$ which takes $1$ on $x$ and 00 elsewhere; and $\langle x\rvert$ for the rank-one operator $\ell^{2}(\Lambda)\to\mathbb{C}$ defined by $\lvert y\rangle\mapsto\left\langle x,y\right\rangle$. Let $T\in\mathbb{B}(\ell^{2}(\Lambda))$, then by $\left\langle x\,\middle|T\middle|\,\mathopen{}y\right\rangle$ we shall mean the inner product of $\left\lvert x\right\rangle$ and $T\left\lvert y\right\rangle$. The collection $(T_{x,y})_{x,y\in\Lambda}$ of $T_{x,y}\mathrel{\vcentcolon=}\langle x\,|T|\,\mathopen{}y\rangle$ are called the matrix elements of $T$.
@@
Let $A$ and $B$ be $\mathbb{Z}/2$-graded “real”C*-algebras. We will write $A\otimes B$ for their graded, minimal tensor product. The grading will cause little difference in this article: we will mostly consider $\mathbb{Z}/2$-graded C*-algebras of the form $A\otimes\mathbb{C}\ell_{p,q}$ or $A\otimes\mathrm{C}\ell_{p,q}$, where $A$ is trivially graded. Then the graded tensor product agrees with the ungraded version. We also write $A\odot B$ for the algebraic tensor product of $A$ and $B$.
Groupoids
A groupoid will sometimes be written as $\mathcal{G}\rightrightarrows\mathcal{G}^{0}$ where $\mathcal{G}^{0}$ is its unit space. If $X,Y\subseteq\mathcal{G}^{0}$, then we write $\mathcal{G}_{X}\mathrel{\vcentcolon=}s^{-1}(X)$, $\mathcal{G}^{Y}\mathrel{\vcentcolon=}r^{-1}(Y)$ and $\mathcal{G}_{X}^{Y}\mathrel{\vcentcolon=}\mathcal{G}_{X}\cap\mathcal{G}^{Y}$. We write $\mathrm{C}^{*}(\mathcal{G})$ for the reduced groupoid C*-algebra of $\mathcal{G}$.
Roe C*-algebras
@@
We write $\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X})$ for the Roe C*-algebra defined by an ample module $(X,\mathcal{H}_{X})$. This differs from the usual convention (e.g. in [Willett-Yu:Higher_index_theory]), and agrees with the convention in [Ewert-Meyer:Coarse_geometry]. In particular, we highlight that $\mathrm{C}^{*}(\mathbb{Z}^{d})$ refers to the group C*-algebra of the group $\mathbb{Z}^{d}$, whereas $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$ refers to the Roe C*-algebra of the discrete metric space $\mathbb{Z}^{d}$.
2. Topological phases and their numerical invariants
When we consider physical systems with anti-unitary symmetries, like time-reversal or particle-hole symmetries, then we must represent them as anti-unitary operators on a Hilbert space which commute or anti-commute with the Hamiltonian. This turns the Hilbert space into a “real” Hilbert space, namely, a Hilbert space that carries an involutive, conjugate-linear map. A concrete C*-algebra on this “real” Hilbert space thus becomes a “real” C*-algebra, that is, a C*-algebra together with a involutive, conjugate-linear *-automorphism, called a “real” structure. Kellendonk [Kellendonk:Cst-algebraic_topological_phases] has shown that the topological phases of such a system are classified by a van Daele’s K-theory [Van_Daele:K-theory_1, Van_Daele:K-theory_2], a model of K-theory that works for graded and “real” Banach algebras, of the corresponding observable “real” C*-algebra. We remark that van Daele’s K-theory coincides with Kasparov theory of “real” C*-algebras, cf. [Roe:Pascheke_duality, Bourne-Kellendonk-Rennie:Cayley_transform, Joseph-Meyer:Geometric_van_Daele_K-theory]. Thus we do not have to distinguish that from the “usual” K-theory if we define the latter using Kasparov theory.
The K-theory $\operatorname{KR}_{*}(A,\mathfrak{r})$ of a “real” C*-algebra $(A,\mathfrak{r})$ is by definition the KO-theory $\operatorname{KO}_{*}(A^{\mathfrak{r}})$ of the real C*-algebra $A^{\mathfrak{r}}$, the fixed-point algebra of the “real” structure $\mathfrak{r}$. In our context, all “real” C*-algebras have simple “real” structures, allowing us to focus on their real subalgebras, and we will simply write $\operatorname{K}_{*}(A)$ for their KO-theory.
Similar to complex C*-algebras, real C*-algebras have a bivariant theory, referred to as KKO-theory; we shall write KK for simplicity when the C*-algebra being real is clear from the context. The difference is that it has a different Bott periodicity. That means, taking suspensions or tensoring with Clifford factors on different entries of this bivariant theory usually yield different groups. Thus we will avoid using $\operatorname{KK}_{n}(A,B)$ and only write $\operatorname{KK}(A\otimes\mathrm{C}\ell_{n,0},B)$ or $\operatorname{KK}(A,B\otimes\mathrm{C}\ell_{n,0})$.
This bivariant theory allows us to describe the index pairings in a formal way. Namely, via the Kasparov product
$\operatorname{KK}(\mathrm{C}\ell_{p,q},A)\times\operatorname{KK}(A\otimes\mathrm{C}\ell_{p^{\prime},q^{\prime}},\mathbb{R})\to\operatorname{KK}(\mathrm{C}\ell_{p+p^{\prime},q+q^{\prime}},\mathbb{R}).$
The range can be identified with a certain real K-theory group of $\mathbb{R}$ by real Bott periodicity.
In our case, the first class in $\operatorname{KK}(\mathrm{C}\ell_{p,q},A)$ comes from the physical Hamiltonian $H$. This is a self-adjoint operator affiliated to $A$, which has a spectral gap at $\mu\in\mathbb{R}$. The symmetries of the system are implemented by the Clifford action. The second class is described by a real spectral triple coming from a tight-binding position operator, acting on the tight-binding Hilbert space $\ell^{2}(\Lambda,\mathcal{K})$. In particular, the assumptions for a spectral triple guarantees that any “physically reasonable” observable C*-algebra $A$ must sit inside the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$, as will be shown in Theorem 3.23.

### v2

@@ 1. Introduction
We begin with some background around it. Topological insulators are materials that are insulating in the bulk but nevertheless possess metallic edges. The current flowing on the edge is usually quite robust under disorder coming from impurity of the crystal. Starting from Bellissard, van Elst and Schulz-Baldes [Bellissard-vElst-SBaldes:NCG_of_QHE, Bellissard:K-theory_in_solid_state_physics], the non-commutative and C*-algebraic tools have been intensively applied to the study of such materials and in particular, leading to an interpretation of the Kubo formula in the integer quantum Hall effect (IQHE) as a quantised index pairing.
@@
These suggest to use more general point sets in $\mathbb{R}^{d}$, modelling the atomic sites of this material, as their underlying geometric spaces. The dynamics thereon are studied by more general actions other than a $\mathbb{Z}^{d}$-translation. It was explained in [Bellissard-Herrmann-Zarrouati:Hulls_aperiodic_solids] that such point sets should be Delone sets (Definition˜3.1). Let $\Lambda$ be a Delone set, whose points are considered as the sites in a physical system. How should one model the observable C*-algebra $A$ from it?
The choice of an observable C*-algebra should be aligned with the following principle: it should be large enough to contain all possible Hamiltonians, and small enough to supply useful homotopy theory (K-theory). In recent years, there have been two main approaches to this modelling problem, which provide toolkits to compute invariants of topological phases:
@@
  • “coarse-geometric” approach describes a C*-algebra which is stable under all short-range, locally–finite-rank perturbations that do not close the gap. This leads to a Roe C*-algebra.
The groupoid approach gives an étale groupoid $\mathcal{G}_{\Lambda}\rightrightarrows\Omega_{0}$ from a Delone set $\Lambda$, and yields a “tight-binding” groupoid C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ , cf. [Bellissard-Herrmann-Zarrouati:Hulls_aperiodic_solids, Bourne-Prodan:Noncommutative_Chern_numbers, Bourne-Mesland:Topological_phases, Mesland-Prodan:Interacting]. This is a generalisation of the well-studied periodic model: if $\Lambda=\mathbb{Z}^{d}$, then $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\simeq\mathrm{C}^{*}(\mathbb{Z}^{d})$ is the group C*-algebra of $\mathbb{Z}^{d}$. Here tight-binding means the following: the regular representation of $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ is given by a Hilbert $\mathrm{C}(\Omega_{0})$-module, or equivalently, a continuous field of Hilbert spaces over $\Omega_{0}$, where $\Omega_{0}$ is unit space of the groupoid $\mathcal{G}_{\Lambda}$. The fibres of this continuous field are canonically identified with $\ell^{2}(\omega)$, where $\omega\in\Omega_{0}$ is a Delone set, either as a translated copy of $\Lambda$, or as a weak*-limit of such sets. Thus $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ consists of families of model Hamiltonians $(H_{\omega})_{\omega\in\Omega_{0}}$, where $H_{\omega}$ acts on $\ell^{2}(\omega)$, in a covariant way with respect to the groupoid action.
The coarse-geometric approach describes a C*-algebra which is stable under all possible short-range perturbations. This gives rise to a (uniform or non-uniform) Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ from $\Lambda$, cf. [Kubota:Controlled_topological_phases, Ewert-Meyer:Coarse_geometry]. We choose to work with the non-uniform Roe C*-algebras, whose advantages over the uniform ones were explained in [Ewert-Meyer:Coarse_geometry].
@@
If we model a topological insulator on a Delone set $\Lambda$ by the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$, then its topological phase is robust in a very strong sense: such topological phases (and their associated topological invariants) are robust under any short-range, locally–finite-rank perturbation that preserves the spectral gap and the symmetry of the system. It follows from the K-theory of (real) Roe C*-algebra (cf. (3.24)) that such topological phase are completely classified by their topological invariants, which are $\mathbb{Z}$- or $\mathbb{Z}/2$-valued indiced that belong to the real K-theory groups of $\mathbb{R}$. The groupoid model $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, on the other hand, has more involved K-theory and numerical invariants. One natural question is in which sense those topological phases described by the groupoid model are still robust. Moreover, do they lead to interesting invariants that do not occur in the coarse-geometric model?
@@
Questions related to this type of robustness are usually answered by showing that the numerical index is still continuous, even for some more general perturbations of the Hamiltonian (which does not have to belong to the observable C*-algebra), together with a “quantisation” statement that the range of the numerical index is $\mathbb{Z}$. This implies that such perturbations leave the topological invariants unchanged.
@@
We describe a different approach to this robustness question by comparing the groupoid model with the coarse-geometric model on $\Lambda$ using the regular representation. The topological invariants defined in [Bourne-Mesland:Topological_phases] factor through this representation. Therefore, since $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ is “robust”, one only needs to understand which topological phases in $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ may still survive there. To this end, we introduce a handy definition of position spectral triples (Definition˜2.11). Both the groupoid model and the coarse-geometric model possess such type of spectral triples. In particular, the position spectral triple $\xi^{\mathrm{Roe}}_{\omega}$ over the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ induces an isomorphism $\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\omega))\to\mathbb{Z}$ (Theorem˜3.33). Then we show that those topological invariants of the groupoid model, given by localising the “bulk cycle” $\prescript{}{d}{\lambda}_{\Omega_{0}}$ at $\omega\in\Omega_{0}$, are pullbacks of the KK-class of $\xi^{\mathrm{Roe}}_{\omega}$ under the comparison *-homomorphism $\mathbb{M}_{N}(\mathrm{C}^{*}(\mathcal{G}_{\Lambda}))\to\mathrm{C}^{*}_{\mathrm{Roe}}(\omega)$ (Theorem˜4.1). Such invariants are therefore strong in the sense of [Ewert-Meyer:Coarse_geometry].
@@
As another application, we show in Theorem˜4.10 that certain numerical invariants on Delone sets, coming from “stacking” lower-dimensional topological phases, must be weak, i.e. unstable under perturbation. This generalises a result in [Ewert-Meyer:Coarse_geometry], which explains why certain topological phases of the periodic model ought to be weak. More precisely, the authors of [Ewert-Meyer:Coarse_geometry] compare the periodic model and the coarse-geometric model using an injective *-homomorphism $\mathrm{C}^{*}(\mathbb{Z}^{d})\hookrightarrow\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$. It vanishes on all but one $\mathbb{Z}$-component of K-theory. In particular, topological phases coming from “stacking” lower-dimensional topological phases will all vanish in $\operatorname{K}_{*}(\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))$. This allows us to conclude that the induced maps $\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d}))\to\operatorname{K}_{*}(\mathrm{C}^{*}(\mathbb{Z}^{d+1}))$ also vanishes there. We show that one can replace $\mathbb{Z}^{d+1}$ by a product Delone set of the form $\Lambda\times L$, where $\Lambda\subseteq\mathbb{R}^{p}$ and $L\subseteq\mathbb{R}^{q}$ are both Delone sets. Using the same strategy in [Ewert-Meyer:Coarse_geometry], we show that stacked topological phases factor through the Roe C*-algebra of a flasque space, which has vanishing K-theory. This allows us to conclude that such topological phases and their numerical invariants have to vanish in the target Roe C*-algebra.
@@ 1.1. Notation and conventions
We fix the following symbols and conventions.
@@
Let $\Lambda$ be a discrete set. We write $\lvert x\rangle$ for the function in $\ell^{2}(\Lambda)$ which takes $1$ on $x$ and 0 elsewhere; and $\langle x\rvert$ for the rank-one operator $\ell^{2}(\Lambda)\to\mathbb{C}$ defined by $\lvert y\rangle\mapsto\left\langle x,y\right\rangle$. Let $T\in\mathbb{B}(\ell^{2}(\Lambda))$, then by $\left\langle x\,\middle|T\middle|\,\mathopen{}y\right\rangle$ we shall mean the inner product of $\left\lvert x\right\rangle$ and $T\left\lvert y\right\rangle$. The collection $(T_{x,y})_{x,y\in\Lambda}$ of $T_{x,y}\mathrel{\vcentcolon=}\langle x\,|T|\,\mathopen{}y\rangle$ are called the matrix elements of $T$.
@@
Let $A$ and $B$ be $\mathbb{Z}/2$-graded “real”C*-algebras. We will write $A\otimes B$ for their graded, minimal tensor product. The grading will cause little difference in this article: we will mostly consider $\mathbb{Z}/2$-graded C*-algebras of the form $A\otimes\mathbb{C}\ell_{p,q}$ or $A\otimes\mathrm{C}\ell_{p,q}$ for a trivial graded (real or complex) C*-algebra $A$. In such cases, the graded tensor product agrees with the ungraded version.
Group(oid) and Roe C*-algebras
@@
We write $\mathrm{C}^{*}(\mathcal{G})$ for the reduced C*-algebra of a groupoid or a group, and $\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X})$ for the Roe C*-algebra defined by an ample module $(X,\mathcal{H}_{X})$. This is aligned with the convention in [Ewert-Meyer:Coarse_geometry]. As a potentially confusing point, we note that $\mathrm{C}^{*}(\mathbb{Z}^{d})$ is the group C*-algebra of the discrete group $\mathbb{Z}^{d}$, and $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$ is the Roe C*-algebra of the discrete metric space $\mathbb{Z}^{d}$.
2. Topological phases and K-theory
In the simplest form, a (non-interacting) quantum system is described by an unbounded, self-adjoint operator $H$ (the Hamiltonian) acting on a complex, separable Hilbert space $\mathcal{H}$. A topological insulator is a quantum system which carries certain symmetries, giving rise to topological invariants that distinguish different topological phases of this system.
The links between topological insulators and (real, bivariant) K-theory has been investigated throughly in the recent decade, cf. [Thiang:K-theoretic_classification, Kellendonk:Cst-algebraic_topological_phases, Bourne-Carey-Rennie:NC_framework_topological_insulators]. In our case of interest, the links are made as follows:
  1. (1)
[…]
  2. (2)
    The topological invariant of a topological insulator can be computed as an index pairing. That is, a Kasparov product of the form
    $\operatorname{KK}(\mathrm{C}\ell_{n,0},A)\times\operatorname{KK}(A\otimes\mathrm{C}\ell_{0,d},\mathbb{R})\to\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R})\simeq\operatorname{KO}_{n-d}(\mathbb{R}),$
    where
    $\operatorname{KO}_{n-d}(\mathbb{R})\simeq\begin{cases}\mathbb{Z}&\text{if }n-d\equiv 0,4{\hphantom{,6,7}}\!\mod 8;\\ \mathbb{Z}/2&\text{if }n-d\equiv 1,2{\hphantom{,6,7}}\!\mod 8;\\ 0&\text{if }n-d\equiv 3,5,6,7\!\mod 8,\end{cases}$
    is the receptacle of $\mathbb{Z}$- or $\mathbb{Z}/2$-valued indices that can be measured in physical experiments.
The section is organised as follows. We shall first recall in Section˜2.1 the definitions of “real” and real C*-algebras and their representations. Then we study two special cases: the Clifford algebras in Section˜2.1.1 and the graded “real” C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$ in Section˜2.1.2. We recall a few properties of Kasparov theory in Section˜2.2 that will be used later.
We introduce a class of spectral triples, called position spectral triples, in Section˜2.3. We recall in Section˜2.3.1 that the position spectral triple $\xi_{\mathbb{Z}^{d}}$ over $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{0,d}$ represents the fundamental class. Via the Kasparov product, it generates a surjective group homomorphism
$\xi]\colon\operatorname{KK}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{0,d})\to\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$